OpenStudy (anonymous):

the width of a rectangle is 7 cm shorter than its length and its area is 228 sq. cm. Find then dimensions of the rectangle?

OpenStudy (anonymous):

w = l - 7 lw = 228 w = 228/l 228 = l^2 - 7l l^2 - 7l + 228 = 0 solve for l, and then use that to find w